OpenStudy (anonymous):

Samuel has a balance on his credit card of $766. If the APR on the credit card is 20.43% compounded monthly, how much is he charged in interest this month?

OpenStudy (anonymous):

$13.04 $20.43 $37.49 $63.83

jimthompson5910 (jim_thompson5910):

I = Prt I = 766(0.2043)*(1/12) I = ???

OpenStudy (anonymous):

Thank you!

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

So if it was compounded daily, the fraction would be 1/365?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

For instance, now I have a reverse question.. Sydney possesses a credit card with an APR of 21.34% compounded daily. What is the APY on this credit card?

jimthompson5910 (jim_thompson5910):

APY = (1 + i/n)^n - 1 APY = (1 + 0.2134/365)^365 - 1 APY = (1 + 0.00058465753424)^365 - 1 APY = (1.00058465753424)^365 - 1 APY = 1.23780251399899 - 1 APY = 0.23780251399899 So APY is roughly 23.78%

jimthompson5910 (jim_thompson5910):

This is effectively the real yearly interest rate

OpenStudy (anonymous):

Okay, I get it. Thank you! I just didnt knwo the formulas

jimthompson5910 (jim_thompson5910):

neither did I, so I had to look it up I found the formula for APY here http://en.wikipedia.org/wiki/APY

jimthompson5910 (jim_thompson5910):

it's not too bad of a formula (compared to others)

OpenStudy (anonymous):

Yeah pretty simple. Thank you!